Every night Scott empties his pocket change into a bowl on his dresser. By Saturday he has 47 coins totaling $5.02. He notices that the number of pennies is the same as the number of quarters and that the sum of the number of pennies and quarters is one more than the sum of the number of nickels and dimes. Write a system of equations to model this problem. Do not solve.
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

p + 5n + 10d + 25q = 502
p = q
p + q = n + d + 1
q = n + d + 1

Answer by ikleyn(52810)

.

            In the post by  @Penguin,  the system of equations is presented  INCORRECTLY.

            I came to bring the correct solution


p + n + d + q = 47
p + 5n + 10d + 25q = 502
p = q
p + q = n + d + 1


-----------


Solved.